Linear Algebra (Cont’d) 3. Let A = 1-X(XTX)-XT. (a) Must A be square? (b) Must X X be square? (e) Must X be square?
Answer #1

3. Given A = I - X(X^TX)^{-1}X^T . We put the respective dimensions aslo, so as to clear up the equation, ie A = \underset{m*m}{I} - \underset{m*n}{X}(\underset{n*m}{X}^T\underset{m*n}{X})^{-1}\underset{n*m}{X}^T . Supposing m and n be positive integers, the identitty matrix will necessarily have equal rows and columns, and supposing X is not a square matrix, having dimesntions m*n, the transpose of X would heve the opposite dimensions, ie n*m.

(a) If X was squared, then A would obviously be squared. Supposing X being a non-square matrix, we have A = \underset{m*m}{I} - \underset{m*n}{X}(\underset{n*m}{X}^T\underset{m*n}{X})^{-1}\underset{n*m}{X}^T or A = \underset{m*m}{I} - \underset{m*n}{X}\underset{n*n}{(X^TX)}^{-1}\underset{n*m}{X}^T (as matrix multiplication of n*m and m*n matrix would produce n*n matrix) or A = \underset{m*m}{I} - \underset{m*m}{X(X^TX)^{-1}X^T} (as matrix multiplication of m*n and n*n and n*m would produce matrix of m*m). Hence, matrix A must be squared, even if m is not equal to n, ie even if X is non-squared.

Note: Inverse is only possible for squared matrix, and the dimesions stays the same after the inversion.

(b) If X was squared, then X^TX would be squared as well. But if X have dimesntions m*n, where m is not equal to n, then X^T's dimension would be the opposite, ie n*m. Hence, (\underset{n*m}{X}^T\underset{m*n}{X}) = \underset{n*n}{(X^TX)} , ie even if X is not squared, X^TX must necessarily be squared.

(c) From above calculations, X must not be necessarily squared to satisfy equation A = I - X(X^TX)^{-1}X^T .